You place 740. g of ice at its melting point in a thermally insulated container. What is the mass of steam at 100°C that would need to be mixed with the ice to produce liquid water at 30.0°C? Any constants you may need can be found in the textbook.
Answer #1

Mass of ice = m1 = 740 g = 0.74 kg

Mass of steam = m2

Initial temperature of ice = T1 = 0 oC

Initial temperature of steam = T2 = 100 oC

Final temperature of the mixture = T3 = 30 oC

Specific heat of water = C = 4186 J/(kg.oC)

Latent heat of fusion of ice = Li = 334000 J/kg

Latent heat of vaporization of steam = Ls = 2260000 J/kg

The heat gained by the ice is equal to the heat lost by the steam.

m1Li + m1C(T3 - T1) = m2Ls + m2C(T2 - T3)

(0.74)(334000) + (0.74)(4186)(30 - 0) = m2(2260000 + 4186(100 - 30))

340089.2 = 2553020m2

m2 = 0.1332 kg

m2 = 133.2 g

Mass of steam needed = 133.2 g